OpenStudy (anonymous):

ps1a version. Hi, below is the version of answer to assignment 1. Interesting to know your approach if it differs

OpenStudy (anonymous):

# find 100th primer primes=[2] #list of primes, staring from 2 qnum=3 #first number we put to 'question' while (len(primes))<=1000 : for jj in range(0,len(primes)): bb = qnum % primes[jj] # generate remainder if bb == 0: # if this remaider is zero ... qnum=qnum+2 # ... then we can thow 'questioned' number away else: # if qnum is not zero we start questioning this 'candidat' from the start for kk in range (0,len(primes)): # and the point is not to miss divisors zz = qnum % primes[kk] if zz == 0: qnum=qnum+2 #print qnum primes.append(qnum) # add 'survived' number to list of primers print "The 10th prime number is" , primes[9] print "The 100th prime number is" , primes[99] print "The 1000th prime number is" , primes[999]

OpenStudy (anonymous):

pls use a code pasting site - http://dpaste.com/667060/ sure takes a long time ... here is mine - http://dpaste.com/667067/

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

http://dpaste.com/hold/667068/

OpenStudy (anonymous):

Your code is a lot faster indeed

OpenStudy (anonymous):

Would you please explain idea behind limit = int(candidate**.5)+2

OpenStudy (anonymous):

i found somewhere that you only have to use divisors up to the the square root of the number that is being tested.

OpenStudy (anonymous):

you can ask in math group to prove it for you :P

OpenStudy (anonymous):

Yes you only need to check up to sqrt(x). http://en.wikipedia.org/wiki/Primality_test "The simplest primality test is as follows: Given an input number n, check whether any integer m from 2 to n − 1 divides n. If n is divisible by any m then n is composite, otherwise it is prime. However, rather than testing all m up to n − 1, it is only necessary to test m up to \scriptstyle\sqrt n"

OpenStudy (anonymous):

I know I may have overkilled it, but here's mine: http://dpaste.com/hold/667157/

OpenStudy (anonymous):

kimci, interesting solution, I reckon, it is fast, easy to understand and straightforward without mathematical tricks

OpenStudy (anonymous):

Thank you all for your suggestions

OpenStudy (anonymous):

Thanks! I haven't tried other bwCA's code, but I think his/hers might be a bit more optimized with the mathematical assumptions and python usage. This is my first time using Python so the syntax I use is what is used at the current lecture I'm at. Which I think I'm only a lecture ahead of Problem sets.